OpenStudy (anonymous):

What is the sum of the finite arithmetic series? (–5) + 0 + 5 + 10 + ... + 65 900 455 450 445

OpenStudy (anonymous):

(–5) + 0 + 5 + 10 + ... + 65 5(-1 + 0 + 1 + 2 + ... + 13) 5(1 + 2 + 3 + ... + 13) - 5 \(\displaystyle\large \left(5\sum_{n=1}^{13}n\right) - 5\)

OpenStudy (anonymous):

\[5 \sum _{n=1}^{13} n-5=450 \]

OpenStudy (anonymous):

Is this clear, @vivianrush ?